# Event-Grounding Graph: Unified Spatio-Temporal Scene Graph from Robotic Observations

### v1

@@ V-A Building EGG from robotic observations
We construct EGG from a stream of multimodal robotic data, including RGB-D images accompanied by frame-wise timestamps and the robot’s pose. The spatial components $\mathcal{G}^{S}$ are implemented as a 3DSG comprising two layers: rooms and objects. While we only concern objects and rooms in our implementation, the formulation from Sec. IV-A does not prevent the inclusion of other spatial concepts, such as buildings and free space from Hydra [2, 3] or asset objects like [14]. We derive object data from an RGB-D video stream to generate spatial nodes. In our implementation, the time-variant attribute set $\mathcal{V}^{i}$ for each object node includes the varying positions of the objects and the state of operation. The position of each object is retrieved by projecting its segmented point cloud from the depth image. While we do not explicitly capture the state of operation, they can be inferred from the instance-wise descriptions $D_{T}$ of the event edges, which we will discuss in more detail later in this section. For simplicity, we record only the initial and final positions of each object for each recorded event. To enrich the object nodes and differentiate between instances of the same semantic class, we gather views of each object from multiple perspectives and utilize a VLM (GPT4o [15]) to generate concise captions that describe their appearance. Thus, the time-invariant attribute set $\mathcal{I}^{i}$ of each object node consists of its name, semantic class, and caption. The room nodes only have time-invariant attributes, which are their manually labeled names and their positions.
@@ V-B Task-based graph pruning
Next, we provide the LLM agent with a list of object names and captions and event summary descriptions from $\mathcal{G}_{Q}^{(1)}$. The agent then, based on the semantic context of the query, is tasked to extract the objects $S_{Q}$ and events $E_{Q}$ related to the query. If there are both objects and events related to the query, we merge them to identify the objects that are most relevant to the query and expand their history. More precisely, if $S_{Q}\neq\emptyset\wedge E_{Q}\neq\emptyset$, then $S^{*}_{Q}=\{n^{S}_{i}\in S_{Q}\ |\ (\exists n^{E}_{T}\in E_{Q})\ e^{E}_{Ti}\in\mathcal{E}^{E(1)}\}$ and $E^{*}_{Q}=\{n^{E}_{T}\in E_{Q}\ |\ (\exists n^{S}_{i}\in S_{Q})\ e^{E}_{Ti}\in\mathcal{E}^{E(1)}\}$, with $\mathcal{E}^{E(1)}$ being the set of event edges of $\mathcal{G}_{Q}^{(1)}$. E.g., if Q=Q="Where is the mug that I was drinking coffee with in the coffee room yesterday?", $S_{Q}=\{\text{all mugs found in the coffee room yesterday}\}$, and $E_{Q}=\{\text{all events involving a person drinking coffee}\}$. After merging, $S^{*}_{Q}$ will contain only the mugs used for drinking coffee in the coffee room yesterday. In the case that $S_{Q}=\emptyset$ or $E_{Q}=\emptyset$, $S_{Q}$ and $E_{Q}$ are not merged. Thus, we retrieve the subgraph:
@@
Finally, if $S_{Q}\neq\emptyset$, we retrieve the task-relevant subgraph $\mathcal{G}_{Q}$ by expanding the event history of relevant objects and pruning the graph to the relevant time interval $T_{Q}$. Intuitively, we expand the history of the relevant objects to retrieve events involving them at all locations in EGG, since objects could move around and events that happen before the key events in $E_{Q}$ could have a role in determining the answer for $Q$. We then prune the graph to the relevant time interval to reduce the graph size and remove the events that are too far in the past or future.
@@ VI-A Experimental setup
To address this gap, we evaluate our framework using real-world data collected from a Hello Robot Stretch 2 mobile manipulator platform with an Astra 2 RGB-D camera mounted on top. Data collection was conducted in two locations within our department building: a coffee room and an office. Our dataset includes 35 videos, captured by the robot, of an individual interacting with 21 objects across the two distinct locations. Additionally, we provide frame-wise timestamps, object instance segmentation, object identification, and local positions of the robot.
The robot task involves verbalizing its observations of a scene through different forms of scene representation. We manually label each video with ground truth summary captions and object role captions to assess the capability of VLMs in comparison to human judgment. For each event, we also manually provide segmentation masks for each object at their first and last appearances, enabling projection to capture their instance positions and incorporate them into the position history of their respective spatial nodes (refer to Sec. V-A for implementation details of the spatial graph). The dataset comprises 80 questions with ground truth answers across four modalities [^2] The full dataset, including the events and question-answer pairs, will be made publicly available after acceptance.:
  • text (24 queries): Returns an answer in natural language. E.g., Describe the mug the person most frequently makes coffee with?
@@
  • node (19): Returns a list of spatial nodes. E.g., Where can I get some coffee?
  • time (10): Returns a time interval or time point. E.g., What is the earliest hour the person was seen working?
To assess the accuracy of the generated responses, we employ LLM semantic scores, where we employ a LLM to judge each generated answer with a score from 0 to 1, for all modalities ($S^{all}$) and separately for text and time modality ($S^{text}$ and $S^{time}$[20]. This evaluation method was chosen over manual evaluation to mitigate potential human biases in judgment and leverage the established capability of LLMs to provide evaluation scores that consider context accuracy and nuance [20]. For the time modality, we also employ semantic scoring to incorporate the contextual nuances of the query. E.g., if the query is "What is the earliest hour in a day that you see someone drinking coffee?", typical methods of calculating time differences may fail to provide the correct answer. This is because the correct response should identify the earliest hour, not just a timestamp closest to a predefined answer. Additionally, we provide quantifiable accuracy scores for node ($A^{node}$) and binary ($A^{binary}$) modalities.
TABLE I: Evaluating the capability of EGG in verbalizing robot observations. The results are averaged over 5 trials.
Method Pruning Captioning $\mathbf{S}^{\mathbf{all}}\uparrow$ $\mathbf{S}^{\mathbf{text}}\uparrow$ $\mathbf{A}^{\mathbf{binary}}\uparrow$ $\mathbf{A}^{\mathbf{node}}\uparrow$ $\mathbf{S}^{\mathbf{time}}\uparrow$ Tokens (1K) $\downarrow$ Compression $\uparrow$
Spatial-only Ground Truth 0.50 0.34 0.50 0.64 0.65 960 48.24 %
Event-only Ground Truth 0.63 0.52 0.74 0.504 0.82 247 84.82 %
EGG (w.o edges) Ground Truth 0.668 0.506 0.688 0.768 0.81 1165 33.09 %
EGG Ground Truth 0.71 0.53 0.72 0.87 0.79 1599 0.00 %
EGG Ground Truth 0.85 0.77 0.87 0.92 0.83 1126 59.42 %
EGG (passed w.o edges) Ground Truth 0.812 0.69 0.88 0.88 0.79 1030 64.52 %
EGG Auto (unguided) 0.68 0.50 0.82 0.73 0.68 1097 60.73 %
EGG Auto (guided) 0.71 0.58 0.80 0.77 0.64 1107 60.04 %
@@ VI-B Experimental results
The results from Tab. I illustrate several ablation studies on different versions of scene representation against our framework. The results were averaged over 5 different trials. In our experiments, GPT4o [15] is utilized as the LLM agent to handle object captioning, graph pruning, and reasoning, and VideoRefer [16] is utilized for video captioning.
VI-B1 The role of each component of EGG
We assess how the unified spatio-temporal features of the EGG framework can preserve the advantages of both purely spatial and temporal representations, enabling the LLM agent to reason more effectively, especially on instance-based queries. Our construction pipeline remains consistent, with the exception that we keep only either the spatial (Spatial-only) or event components (Event-only), or keep both components and omit the event edges (w.o edges) to create the incomplete versions of EGG. We omit graph pruning for the incomplete versions of EGG since, without event-object linkage, the pruning process would be overly aggressive and potentially remove relevant information. Additionally, to assess the importance of edge instance-wise descriptions, we perform pruning on the original EGG and omit the event edges when serializing the graph to JSON format (passed w.o edges). The results for this study are presented in Tab. I.
Qualitatively, the spatial-only version of EGG, i.e., a 3DSG, struggles to address queries that depend on event-related information. Conversely, the event-only version fails to adequately respond to queries that necessitate knowledge of which object was involved in the event. E.g., for the query $Q_{1}$, "Which mug was used for coffee-making on August 30, 2025?", without grounding coffee-making events to specific object nodes, neither the spatial-only nor the event-only representations can answer correctly. The version that includes both spatial and event components without connecting them with edges (w.o edges) improves performance for questions that involve knowledge of object instances, thanks to the inclusion of spatial nodes, as evident in the improved performance in $A^{node}$. However, similar to the event-only version, it still falters at queries that involve granular interactions, which are absent due to the lack of connections between objects and events. Without connecting events to spatial nodes, information retrieval for queries regarding object types that have multiple instances, e.g., mugs, is impossible because objects cannot be differentiated.
When looking at the results involving pruning, the version that omits edges when serializing performs close to the full version, but misses some needed details for understanding specific interactions. Thanks to the pruning process, the resulting graph removes unrelated objects of the same class, e.g., for the same query $Q_{1}$, only the mugs involved in the coffee-making events are present in the subgraph $\mathcal{G}_{Q_{1}}$. Furthermore, we observe that there are some instances in which the LLM agent can use the overlapping timestamps between the object’s recorded position and the event to infer the relationship between the objects and the events in the absence of event-object edges.
@@ VI-B1 The role of each component of EGG
Overall, it is evident that each component of EGG—the spatial, event components, and the connecting edges between events and objects—plays a crucial role in enriching the information pool provided to the LLM agent. While complete edge information is valuable, the pruned version without edges can still offer reasonable performance if token limits are a consideration.
VI-B2 Graph pruning efficiency
@@
To assess how the subgraph $\mathcal{G}_{Q}$ affects reasoning performance and token usage, we compare two versions of the information search strategy: one utilizing the pruning method detailed in Sec. V-B and the other using the complete graph with the LLM agent (EGG with and without pruning in Tab. I). The version employing graph pruning outperforms the alternative and achieves approximately 60% graph compression, resulting in a 30% reduction in token usage. It should be noted that the token count includes the input prompt tokens. The pruned version of EGG demonstrated significantly higher overall semantic score in the downstream information retrieval task. Disregarding the excess input tokens, we also hypothesize that this improvement may stem from the LLM’s tendency to hallucinate when presented with redundant information, as discussed in studies like [11]. Moreover, LLMs have shown positional biases towards information presented at the beginning or end of the input prompt [21]. Therefore, it is evident that LLMs perform more effectively when provided only with relevant context to the specific task. The experimental results indicate that the pruning process improves LLM overall semantic score on downstream tasks by excluding task-irrelevant entities from the representation.
VI-B3 Video captioning capability
@@
In this study, we construct EGG using three different video captioning methods: using ground truth labeling and using a VLM, VideoRefer [16], with and without guided captioning. Guided captioning utilizes a ground truth list of objects pertinent to the event, followed up by automated captioning. As anticipated, the quality of video captioning directly affects the accuracy of subsequent tasks, as detailed in Tab. I.
Our findings indicate that video captioning remains a bottleneck in achieving human-level performance. Specifically, when EGG is constructed with guided captioning, there is a 14% decrease in overall semantic score compared to using ground truth event labeling. For EGG built with unguided captioning, the performance drop is even more pronounced, at 17%. These results underscore the necessity for advancements in video captioning models.
VI-B4 Failure modes analysis
@@
Fig. 3 provides a summary of the potential failure modes identified during an experimental trial, as detailed in Tab. I. Our observations show that most errors originate from the LLM answer generation. More specifically, the primary sources of error include:
@@
  • Correct TMI (Too much information): while the LLM may provide the correct answer, it often includes excessive, redundant information.
In contrast, we identified only a single instance where pruning removed a vital node related to the query, suggesting that EGG provides adequate information for reasoning tasks and effectively supports graph-based reasoning with current LLMs, and the shortcomings mostly arise from the interaction with the LLM.
@@ VII Discussions
Our experiments demonstrate that EGG’s grounding of observed dynamic events with spatial features allows it to retain the capabilities of purely spatial and temporal representations while improving understanding of historical interactions within a scene. We have also shown that incorporating a graph pruning strategy to select relevant information improves the performance on the downstream information retrieval task and reduces token usage.
However, a notable constraint of our framework is that the subgraph $\mathcal{G}_{Q}$ as formulated in (1) is not guaranteed to be optimal. We observe that often the number of relevant objects typically returned can be redundant, as our heuristic approach may include irrelevant objects connected to pertinent events. This redundancy also applies to the events. Thus, this underscores the necessity for the development of more robust semantic search algorithms. Another constraint of EGG lies in the lack of connectivity between the recorded events. As a result, the LLM agent has to infer from scattered events and form implicit links based on the timestamps or event ordering. A hierarchical structure in the time domain, such as having summaries of multiple events [5], could further improve the performance of our framework. Moreover, the process of building EGG is currently offline, making it not yet suitable for real-time robotic applications. With the advancements of memory-efficient VLMs [22, 4] and the availability of real-time updatable 3DSG frameworks [13], the prospect of deploying a robot to construct a unified spatio-temporal representation in real-time will become increasingly feasible.

### v2

@@ V-A Building EGG from robotic observations
We construct EGG from a stream of multimodal robotic data, including RGB-D images accompanied by frame-wise timestamps and the robot’s pose. We derive object data from an RGB-D video stream to generate spatial nodes. In our implementation, the time-variant attribute set $\mathcal{V}^{i}$ for each object node includes the varying positions of the objects and the state of operation. The position of each object is retrieved by projecting its segmented point cloud from the depth image. While we do not explicitly capture the state of operation, they can be inferred from the instance-wise descriptions $D_{T}$ of the event edges, which we will discuss in more detail later in this section. For simplicity, we record only the initial and final positions of each object for each recorded event. To enrich the object nodes and differentiate between instances of the same semantic class, we gather views of each object from multiple perspectives and utilize a VLM (GPT4o [19]) to generate concise captions that describe their appearance. Thus, the time-invariant attribute set $\mathcal{I}^{i}$ of each object node consists of its name, semantic class, and caption. The room nodes only have time-invariant attributes, which are their manually labeled names and their positions.
@@ V-B Task-based graph pruning
Next, we provide the LLM agent with a list of object names and captions and event summary descriptions from $\mathcal{G}_{Q}^{(1)}$. The agent then, based on the semantic context of the query, is tasked to extract the objects $O_{Q}$ and events $E_{Q}$ related to the query. If there are both objects and events related to the query, we merge them to identify the objects that are most relevant to the query and expand their history. More precisely, if $O_{Q}\neq\emptyset\wedge E_{Q}\neq\emptyset$, then $O^{*}_{Q}=\{n^{S}_{i}\in O_{Q}\ |\ (\exists n^{E}_{T}\in E_{Q})\ e^{E}_{Ti}\in\mathcal{E}^{E(1)}\}$ and $E^{*}_{Q}=\{n^{E}_{T}\in E_{Q}\ |\ (\exists n^{S}_{i}\in O_{Q})\ e^{E}_{Ti}\in\mathcal{E}^{E(1)}\}$, with $\mathcal{E}^{E(1)}$ being the set of event edges of $\mathcal{G}_{Q}^{(1)}$. E.g., if Q=Q="Where is the mug that I was drinking coffee with in the coffee room yesterday?", $O_{Q}=\{\text{all mugs found in the coffee room yesterday}\}$, and $E_{Q}=\{\text{all events involving a person drinking coffee}\}$. After merging, $O^{*}_{Q}$ will contain only the mugs used for drinking coffee in the coffee room yesterday. In the case that $O_{Q}=\emptyset$ or $E_{Q}=\emptyset$, $O_{Q}$ and $E_{Q}$ are not merged. Thus, we retrieve the subgraph:
@@
Finally, if $O_{Q}\neq\emptyset$, we retrieve the task-relevant subgraph $\mathcal{G}_{Q}$ by expanding the event history of relevant objects and pruning the graph to the relevant time interval $T_{Q}$. Intuitively, we expand the history of the relevant objects to retrieve events involving them at all locations in EGG, since objects could move around and events that happen before the key events in $E_{Q}$ could have a role in determining the answer for $Q$. We then prune the graph to the relevant time interval to reduce the graph size and remove the events that are too far in the past or future.
@@ VI-A Experimental setup
To address this gap, we evaluate our framework using real-world data collected from a Hello Robot Stretch 2 mobile manipulator platform with an Astra 2 RGB-D camera mounted on top. Data collection was conducted in two locations within our department building: a coffee room and an office. Our dataset includes 35 videos, captured by the robot, of an individual performing various daily activities and interacting with 21 distinct objects in a coffee room and an office. On average, an event consists of 3 to 4 objects being interacted with. Additionally, we provide frame-wise timestamps, object instance segmentation, event-relevant object identification, and local positions of the robot.
The robot task involves verbalizing its observations of a scene through different forms of scene representation. We manually label each video with ground truth summary captions and object role captions to assess the capability of VLMs in comparison to human judgment. For each event, we also manually provide segmentation masks for each object at their first and last appearances, enabling projection to capture their instance positions and incorporate them into the position history of their respective spatial nodes (refer to Sec. V-A for implementation details of the spatial graph). The dataset comprises 80 questions with ground truth answers across four modalities, for instance [^2] The full dataset, including the events and question-answer pairs, will be made publicly available after acceptance to preserve anonymity.:
  • text (24 queries): Returns an answer in natural language. E.g., Which room was the phone last seen used in?
@@
  • node (18): Returns a list of spatial nodes. E.g., The person left something on the office table at some point between 15:30:00 and 16:16:00. What was it?
  • time (10): Returns a point in time. E.g., What is the earliest time the person was seen working?
To assess the accuracy of the generated responses for the queries with text modality we employ LLM semantic scores, where we employ a LLM to judge each generated answer with a score from 0 to 1. This evaluation method was chosen over manual evaluation to mitigate potential human biases in judgment and leverage the established capability of LLMs to provide evaluation scores that consider context accuracy and nuance [13]. For the time modality, we consider an answer to be correct if it lies within 2 minutes of the ground truth answer. Additionally, we provide quantifiable F1 scores for node ($F1^{node}$) and binary ($F1^{binary}$) modalities.
In our experiments, GPT4o [19] is utilized to handle object captioning, graph pruning, and reasoning, and VideoRefer [31] is utilized for automatic video captioning.
Method $\mathbf{S}^{\mathbf{text}}\uparrow$ $\mathbf{F1}^{\mathbf{binary}}\uparrow$ $\mathbf{F1}^{\mathbf{node}}\uparrow$ $\mathbf{A}^{\mathbf{time}}\uparrow$
EGG 0.70 0.78 0.76 0.78
Hydra* [11, 12] 0.52 0.00 0.68 0.10
ReMEmbR [1] 0.54 0.49 0.38
    Implemented by removing the event nodes $\mathcal{N}^{E}$ and event edges $\mathcal{E}^{E}$ from EGG
  • Since ReMEmbR does not have access to the scene graph, we did not evaluate it on the queries with node modality.
TABLE I: Evaluating the capability of EGG in verbalizing robot observations against baseline methods. The results are averaged over 5 trials. VideoRefer [31] is used for automatic captioning of events for EGG and ReMEmbR.
@@ VI-B Experimental results
We evaluate our method against two baselines, a purely spatial memory representation, Hydra [11, 12], and a spatio-temporal memory building method, ReMEmbR [1]. Our experimental results, presented in Tab. I, clearly demonstrate that our method outperforms both baselines across all modalities. From these findings, it can be inferred that the unified spatio-temporal features of EGG preserve the advantages of both purely spatial and temporal representations, enabling the LLM agent to reason more effectively, especially on instance-based queries.
Qualitatively, Hydra [11, 12], a purely spatial scene representation, struggles to address event-related queries. By observing the reasoning output, the LLM often rely solely on the positions of objects to infer their relationships. However, this approach is ineffective for queries that require detailed information about human interactions, which are not typically captured in a 3DSG. For instance, the query "Describe the appearance of the most frequently used mug for making coffee at the coffee machine" requires usage frequency of different mugs to correctly answer.
On the other hand, ReMEmbR [1] performs better on descriptive queries with its access to event data. Yet, it lacks the grounding of events with spatial features, so the LLM can only rely on video captions. Notably, it fails to address queries such as "How many different mugs were seen in the events?", which could be easily answered with spatial representations. Furthermore, it struggles with queries requiring the usage history of specific objects, such as "Describe the appearance of the most frequently used mug for making coffee at the coffee machine".
@@
Overall, it is evident that each component of EGG—the spatial, event components—plays a crucial role in enriching the information provided to the LLM agent.
VI-C Ablation study
TABLE II: Ablation study on EGG. The results are averaged over 5 trials.*
Method $\mathbf{A}^{\mathbf{all}}\uparrow$ $\mathbf{S}^{\mathbf{text}}\uparrow$ $\mathbf{F1}^{\mathbf{binary}}\uparrow$ $\mathbf{F1}^{\mathbf{node}}\uparrow$ $\mathbf{S}^{\mathbf{time}}\uparrow$ Tokens (1K) $\downarrow$ Compression $\uparrow$
EGG 0.83 0.85 0.77 0.87 0.80 998 64.72 %
EGG (w.o edges) 0.79 0.81 0.72 0.82 0.80 922 68.02 %
EGG (w.o pruning) 0.77 0.80 0.65 0.79 0.88 1576 0.00 %
EGG (auto-captions) 0.76 0.70 0.78 0.76 0.78 985 66.55 %
    Ground truth captioning is used for event nodes in all versions except for auto-captions, for which VideoRefer [31] is used.
The results from Tab. II illustrate several ablation studies on different versions of EGG. In addition to the metrics mentioned in Sec. VI-A, we provide an additional accuracy score across all modalities ($A^{all}$), a tokens count and graph compression percentage, which is the ratio of the number of characters of the JSON-serialized subgraph $\mathcal{G}_{Q}$ with the JSON-serialized full graph $\mathcal{G}$.
VI-C1 The role of event edges
To assess the importance of event edges, we perform pruning on the original EGG and omit them when JSON-serializing the graph (w.o edges in Tab. II). This version of EGG performs close to the standard version, but misses some needed details for understanding specific interactions. In most cases, the pruning process essentially removes unrelated objects of the same class from the resulting graph. E.g., for the query $Q_{1}=$"Which mug was most often used for coffee-making?", only the mugs involved in the coffee-making events are present in the subgraph $\mathcal{G}_{Q_{1}}$. Although complete edge information is valuable and the standard version shows higher performance, removing edges after pruning can still provide reasonable performance when token limits are a consideration.
VI-C2 Graph pruning efficiency
@@
To assess how the subgraph $\mathcal{G}_{Q}$ affects reasoning performance and token usage, we compare two versions of the information search strategy: one utilizing the pruning method detailed in Sec. V-B and the other using the complete graph with the LLM agent (EGG w.o pruning in Tab. II). The version employing graph pruning outperforms the alternative and achieves approximately 65% graph compression, resulting in a 36% reduction in token usage. It should be noted that the token count includes the input prompt tokens. The pruned version of EGG demonstrated significantly higher overall semantic score in the downstream information retrieval task. Disregarding the excess input tokens, we also hypothesize that this improvement may stem from the LLM’s tendency to hallucinate when presented with redundant information, as discussed in studies like [26]. Moreover, LLMs have shown positional biases towards information presented at the beginning or end of the input prompt [28]. The experimental results indicate that the pruning process improves LLM performance on downstream tasks by excluding task-irrelevant entities from the representation.
VI-C3 Video captioning capability
@@
In this study, we construct EGG using two different video captioning methods: using ground truth labeling and using a VLM, VideoRefer [31]. As anticipated, the quality of video captioning directly affects the accuracy of subsequent tasks. As detailed in Tab. I, automatic captioning reduces the overall accuracy by 9% compared to using ground truth event labeling. These results underscore the necessity for advancements in video captioning models.
VI-D Failure modes analysis
@@
Fig. 3 provides a summary of the potential failure modes identified during an experimental trial using ground truth event captions (shown in Tab. II). Our observations show that most errors originate from the LLM answer generation. More specifically, the primary sources of error include:
@@
In contrast, we identified only 1 instance where pruning removed a vital node related to the query, suggesting that EGG provides adequate information for reasoning tasks and effectively supports graph-based reasoning with current LLMs, and the shortcomings mostly arise from the interaction with the LLM.
@@ VII Discussions
Our experiments demonstrate that EGG’s grounding of observed dynamic events with spatial features allows it to retain the capabilities of purely spatial and temporal representations while improving understanding of historical interactions within a scene. We have also shown that incorporating a graph pruning strategy to select relevant information improves the performance on the downstream information retrieval task and reduces token usage. By retaining both geometric changes and semantic context, EGG could provide valuable information for robots to execute higher level tasks, such as making more informed decisions during semantic object search [7, 9], or learning mobile manipulation skills from human demonstrations [16].
However, a notable constraint of our framework is that the subgraph $\mathcal{G}_{Q}$ as formulated in (1) is not guaranteed to be optimal. We observe that often the number of relevant objects typically returned can be redundant, as our heuristic approach may include irrelevant objects connected to pertinent events. This redundancy also applies to the events. Thus, this underscores the necessity for the development of more robust semantic search algorithms, especially when extending to larger, more complex environment with more events. Another constraint of EGG lies in the lack of connectivity between the recorded events. As a result, the LLM agent has to infer from scattered events and form implicit links based on the timestamps or event ordering. A hierarchical structure in the time domain, such as having summaries of multiple events [3], could further improve the performance of our framework. Moreover, the process of building EGG is currently offline, making it not yet suitable for real-time robotic applications. Currently, our method requires that any information queried be already present in the map, a limitation shared by many semantic mapping techniques. This issue is however further compounded by the addition of event memory. Efforts to overcome this limitation, such as those that focus on inferring unseen interactions and transitions [24], are crucial and hold great promise for long-term robotic operation.